student was asked to form two-digit numbers using the following digits: 0,1,2,3,4,5,6,7,
and 8
how many of these are divisable by 2?
Repeating of digits are not allowed
Can you help me with this?.... thank you so much

Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)


Hi
form two-digit numbers using the following
9 numbers: 0,1,2,3,4,5,6,7,8
how many of these are divisible by 2? Even
tens digit 8 choices (not zero)
ones digit 5 choices (0,2,4,6,8)
8x5 = 40 choices for even numbers
Repeating of digits are not allowed (22,44,66,88)
40 - 4 = 36 is the number of even with no repeating digits
Wish You the Best in your Studies.

Answer by ikleyn(52781)

.

First digit can be any of the 8 given digits from 1 to 8 inclusive.

    (notice that 0 (zero) can not be the first digit of a two-digit number).
 


Second digit can be any of the 5 given EVEN digits 0, 2, 4, 6, and 8.



It gives us  8*5 = 40 potential candidates for two digit numbers.



From this amount, we should exclude two-digit numbers with repeating digits.

They are  22, 44, 66, and 88 - in all, 4 numbers must be excluded from 40.


It gives the ANSWER  40 - 4 = 36.


ANSWER.  There are 36 two-digit numbers satisfying to imposed conditions.

Solved.